Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $1.00 coming 3 years from today. The dividend should grow rapidly - at a rate of 19% per year - during Years 4 and 5, but after Year 5, growth should be a constant 5% per year. If the required return on Computech is 15%, what is the value of the stock today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Year 3 dividend = 1

Year 4 dividend = 1 (1 + 19%) = 1.19

Year 5 dividend = 1.9 (1 + 19%) = 1.4161

Year 6 dividend = 1.4161 (1 + 5%) = 1.486905

Value at year 5 = D6 / required rate - growth rate

Value at year 5 = 1.486905 / 0.15 - 0.05

Value at year 5 = 1.486905 / 0.1

Value at year 5 = 14.8691

Value of stock = 1 / (1 + 0.15)3 + 1.19 / (1 + 0.15)4 + 1.4161 / (1 + 0.15)5 + 14.8691 / (1 + 0.15)5

Value of stock = $9.43